OpenStudy (anonymous):

v= (3.14) r^2h for h

OpenStudy (anonymous):

is it v = (3.14)* r^(2h) or v = (3.14)* r^(2) * h ????

OpenStudy (anonymous):

v= (3.14)r^(2)* h

OpenStudy (across):

That's the formula for the volume of a cylinder: \[V/\pi r^2=h\]

OpenStudy (anonymous):

indeed haha. h = v/(pi*r^2)